# Deep Reinforcement Learning-Based Dynamic Resource Allocation in Cell-Free Massive MIMO

### v1

@@ I Introduction
Recently, cell-free massive multiple-input multiple-output (CFmMIMO) has emerged as a promising technology for next generation wireless systems. In CFmMIMO networks, a large number of access points (APs) are deployed in a large area to coherently serve users and offer uniformly great quality of service (QoS) to all users with simple signal processing [1, 2]. However, dense AP deployment in CFmMIMO imposes significant energy demands, making energy efficiency (EE) a key priority for the system design [3].
Maximizing EE for CFmMIMO systems often entails a joint optimization of power allocation together with the dynamic selection of APs and/or antennas, both involving high-dimensional variables due to the large size of CFmMIMO networks [2, 4]. Conventional optimization approaches to EE maximization, such as successive convex approximation (SCA) [2] and accelerated projected gradient [5] often offer good performance. However, they are generally iterative in nature, requiring high computational complexity and long run time. Thus, these methods have limited practicality for real-time operation of large-size systems as CFmMIMO [2, 6, 7, 5].
Deep reinforcement learning (DRL) has emerged as a promising paradigm for real-time operations [8]. However, existing DRL frameworks for EE design and optimization in CFmMIMO systems face critical scalability issues. Specifically, DRL approaches for power control [9] and user association [10, 11] often require action spaces whose dimensionality grows with the network size. Similarly, multi-agent DRL schemes [12] typically scale the number of agents with the number of users, further amplifying the scalability issue. Therefore, these DRL solutions can be unsuitable for large-scale CFmMIMO networks. Conversely, scalable DRL designs that randomly deactivate APs according to deployment density oversimplify the problem and can result in suboptimal performance [13]. Therefore, developing a DRL framework that is simultaneously scalable and capable of jointly optimizing power allocation, AP deactivation, and per-AP antenna selection to maximize EE remains a critical open challenge.
@@ II System Model
We consider a CFmMIMO system in which $M$ APs, each equipped with $N$ antennas, serve $K$ single‑antenna users over the same time‑frequency resources, with $MN\gg K$ [2]. All APs connect to a central processing unit (CPU) via a backhaul network, as illustrated in Fig. 1. We consider the scenario that at each AP, only a subset of antennas is activated to reduce the power consumption. Let $N_{m}\leq N$ be the number of activated antennas at AP $m$.
@@ II-A Channel Estimation
We define the effective channel between user $k$ and the active antennas of AP $m$ as ${\mathbf{g}}_{mk}^{\text{A}}\triangleq{\mathbf{s}}_{m}\odot{\mathbf{g}}_{mk}$. The effective channel ${\mathbf{g}}_{mk}^{\text{A}}$ therefore contains only the components corresponding to the active antennas, which are involved in channel estimation and data transmission. To estimate ${\mathbf{g}}_{mk}^{\text{A}}$, ${\mathbf{Y}}_{\mathsf{p},m}$ is projected onto $\bm{\varphi}_{k}$ to yield $\tilde{{\mathbf{y}}}_{\mathsf{p},mk}\triangleq{\mathbf{Y}}_{\mathsf{p},m}\bm{\varphi}_{k}$. The minimum mean-square error (MMSE) estimate of the channel ${\mathbf{g}}_{mk}$ is given by [14]
@@
Let $\hat{g}^{\text{A}}_{mkn}$ denote the $n$-th element of $\hat{\mathbf{g}}^{\text{A}}_{mk}$. Due to antenna selection, we have $\hat{g}^{\text{A}}_{mkn}=0$ if $s_{mn}=0$ , while for $s_{mn}=1$, the mean square of $\hat{g}^{\text{A}}_{mkn}$ is given as [2]
@@ III-A Spectral Efficiency
Each user detects its signal based on statistical channel knowledge, eliminating the need for downlink pilots [1, 15]. The received signal $r_{k}$ in (5) can be decomposed as
@@
Here, $\mathrm{DS}_{k}\triangleq\sqrt{\rho_{\mathsf{d}}}\mathbb{E}\{\sum_{m=1}^{M}\sqrt{\eta_{mk}}{\mathbf{u}}_{mkk}\}$, $\mathrm{BU}_{k}\triangleq\sqrt{\rho_{\mathsf{d}}}(\sum_{m=1}^{M}\sqrt{\eta_{mk}}{\mathbf{u}}_{mkk}-\mathbb{E}\{\sum_{m=1}^{M}\sqrt{\eta_{mk}}{\mathbf{u}}_{mkk}\})$, and $\mathrm{UI}_{kj}\triangleq\sqrt{\rho_{\mathsf{d}}}\sum_{m=1}^{M}\sqrt{\eta_{mj}}{\mathbf{u}}_{mkj}$, with ${\mathbf{u}}_{mkj}\triangleq({\mathbf{g}}_{mk}\odot{\mathbf{s}}_{m})^{\scriptscriptstyle\mathsf{T}}\hat{{\mathbf{g}}}_{mj}^{*}$, represent the desired signal, beamforming uncertainty, and inter-user interference power, respectively. Accordingly, the SE for user $k$ is then given by [15]
@@ IV-A Resource Allocation
Power allocation: We adopt the fractional power allocation method [16], which assigns more power to users with larger values of $\gamma_{mk}$, $m\in\mathcal{A}_{\mathsf{active}}$, where we recall that $\gamma_{mk}$ denotes the estimated channel gain as defined in (3). The power allocated by AP $m$ to user $k$ is then obtained as
@@ IV-B DRL Agent Formulation
With the resource allocation mechanisms described above, antenna selection and power allocation for EE maximization reduce to the joint optimization of $\{\zeta,\kappa,\nu\}$. We formulate this as a reinforcement learning (RL) task [17], where a RL agent adaptively adjusts these parameters based on the observed network states. It is worth noting that all of these parameters depend only on the large-scale fading coefficients $\beta_{mk}$. Consequently, the stage remains unchanged within one large-scale fading interval. Let $t\in\{1,2,\ldots,T\}$ index the large-scale intervals, hereafter referred to as “time slots” for simplicity.
@@ IV-C Proximal Policy Optimization
To solve the RL problem in (16), we employ proximal policy optimization (PPO) [18], which effectively handles continuous variables and ensures stable learning by restricting policy updates to small, controlled steps. In PPO, the agent’s policy $\pi_{\theta}({\mathbf{a}}_{t}|{\mathbf{s}}_{t})$ and value function $V_{\phi}({\mathbf{s}}_{t})$ are parameterized by neural networks with parameters $\theta$ and $\phi$, respectively. The policy network outputs the mean and variance of a continuous action distribution from which ${\mathbf{a}}_{t}$ is sampled, while the value network estimates the expected return $V_{\phi}({\mathbf{s}}_{t})$ to guide policy updates. The advantage function, estimated via generalized advantage estimation (GAE) [18], is given by
@@ V Performance Evaluation
We consider a square network area of $1\times 1$ km2, wherein the APs and users are uniformly and independently distributed, with wrap-around at the edges to avoid boundary effects. The large-scale fading coefficient is modeled as $\beta_{mk}=\mathrm{PL}_{mk}\cdot z_{mk}$, where $z_{mk}$ is log-normal shadowing with standard deviation $\sigma_{\mathrm{sh}}=8$ dB. The path loss $\mathrm{PL}_{mk}$ (in dB) follows the three-slope model [2]:
@@
The DRL agent is trained using the PPO algorithm [18]. Both the actor and critic networks consist of two hidden layers with 256 neurons each and ReLU activations. The learning rate and reward penalty coefficient are set to $3\times 10^{-4}$ and $\xi_{\mathrm{pen}}=20$, respectively, and the agent is trained for 300,000 timesteps using a minibatch size of 64. Other key hyperparameters of the PPO algorithm are the discount factor $\vartheta=0.99$, the clipping parameter $\epsilon=0.2$, and the GAE parameter $\lambda=0.95$.
We compare the proposed DRL scheme against three benchmarks: The DRL-AP scheme assumes all APs are active ($\zeta=1$) and optimizes only antenna allocation ($\kappa$) and power control ($\nu$), isolating the effect of AP deactivation. The DRL-AO scheme optimizes only AP activation ($\zeta$), with $\kappa=0$ and $\nu=1$, highlighting the benefits of network sparsification. Finally, FA-SCA uses all antennas and optimizing power allocation using the SCA method [2].

### v2

@@ I Introduction
Recently, cell-free massive multiple-input multiple-output (CFmMIMO) has emerged as a promising technology for next generation wireless systems. In CFmMIMO networks, a large number of access points (APs) are deployed in a large area to coherently serve users and offer uniformly great quality of service (QoS) to all users with simple signal processing [ngo2017cell, ngo2017total]. However, dense AP deployment in CFmMIMO imposes significant energy demands, making energy efficiency (EE) a key priority for the system design [ngo2024ultradense].
Maximizing EE for CFmMIMO systems often entails a joint optimization of power allocation together with the dynamic selection of APs and/or antennas, both involving high-dimensional variables due to the large size of CFmMIMO networks [ngo2017total, nguyen2022hybrid]. Conventional optimization approaches to EE maximization, such as successive convex approximation (SCA) [ngo2017total] and accelerated projected gradient [mai2022energy] often offer good performance. However, they are generally iterative in nature, requiring high computational complexity and long run time. Thus, these methods have limited practicality for real-time operation of large-size systems as CFmMIMO [ngo2017total, alonzo2019energy, bashar2019energy, mai2022energy].
Deep reinforcement learning (DRL) has emerged as a promising paradigm for real-time operations [wang2024green]. However, existing DRL frameworks for EE design and optimization in CFmMIMO systems face critical scalability issues. Specifically, DRL approaches for power control [luo2022downlink] and user association [ghiasi2022energy, mendoza2023user] often require action spaces whose dimensionality grows with the network size. Similarly, multi-agent DRL schemes [liu2024joint] typically scale the number of agents with the number of users, further amplifying the scalability issue. Therefore, these DRL solutions can be unsuitable for large-scale CFmMIMO networks. Conversely, scalable DRL designs that randomly deactivate APs according to deployment density oversimplify the problem and can result in suboptimal performance [topal2023drl]. Therefore, developing a DRL framework that is simultaneously scalable and capable of jointly optimizing power allocation, AP deactivation, and per-AP antenna selection to maximize EE remains a critical open challenge.
@@ II System Model
We consider a CFmMIMO system in which $M$ APs, each equipped with $N$ antennas, serve $K$ single‑antenna users over the same time‑frequency resources, with $MN\gg K$ [ngo2017total]. All APs connect to a central processing unit (CPU) via a backhaul network, as illustrated in Fig. 1. We consider the scenario that at each AP, only a subset of antennas is activated to reduce the power consumption. Let $N_{m}\leq N$ be the number of activated antennas at AP $m$.
@@ II-A Channel Estimation
We define the effective channel between user $k$ and the active antennas of AP $m$ as ${\mathbf{g}}_{mk}^{\text{A}}\triangleq{\mathbf{s}}_{m}\odot{\mathbf{g}}_{mk}$. The effective channel ${\mathbf{g}}_{mk}^{\text{A}}$ therefore contains only the components corresponding to the active antennas, which are involved in channel estimation and data transmission. To estimate ${\mathbf{g}}_{mk}^{\text{A}}$, ${\mathbf{Y}}_{\mathsf{p},m}$ is projected onto $\bm{\varphi}_{k}$ to yield $\tilde{{\mathbf{y}}}_{\mathsf{p},mk}\triangleq{\mathbf{Y}}_{\mathsf{p},m}\bm{\varphi}_{k}$. The minimum mean-square error (MMSE) estimate of the channel ${\mathbf{g}}_{mk}$ is given by [kay1993fundamentals]
@@
Let $\hat{g}^{\text{A}}_{mkn}$ denote the $n$-th element of $\hat{\mathbf{g}}^{\text{A}}_{mk}$. Due to antenna selection, we have $\hat{g}^{\text{A}}_{mkn}=0$ if $s_{mn}=0$ , while for $s_{mn}=1$, the mean square of $\hat{g}^{\text{A}}_{mkn}$ is given as [ngo2017total]
@@ III-A Spectral Efficiency
Each user detects its signal based on statistical channel knowledge, eliminating the need for downlink pilots [ngo2017cell, marzetta2016fundamentals]. The received signal $r_{k}$ in (5) can be decomposed as
@@
Here, $\mathrm{DS}_{k}\triangleq\sqrt{\rho_{\mathsf{d}}}\mathbb{E}\{\sum_{m=1}^{M}\sqrt{\eta_{mk}}{\mathbf{u}}_{mkk}\}$, $\mathrm{BU}_{k}\triangleq\sqrt{\rho_{\mathsf{d}}}(\sum_{m=1}^{M}\sqrt{\eta_{mk}}{\mathbf{u}}_{mkk}-\mathbb{E}\{\sum_{m=1}^{M}\sqrt{\eta_{mk}}{\mathbf{u}}_{mkk}\})$, and $\mathrm{UI}_{kj}\triangleq\sqrt{\rho_{\mathsf{d}}}\sum_{m=1}^{M}\sqrt{\eta_{mj}}{\mathbf{u}}_{mkj}$, with ${\mathbf{u}}_{mkj}\triangleq({\mathbf{g}}_{mk}\odot{\mathbf{s}}_{m})^{\scriptscriptstyle\mathsf{T}}\hat{{\mathbf{g}}}_{mj}^{*}$, represent the desired signal, beamforming uncertainty, and inter-user interference power, respectively. Accordingly, the SE for user $k$ is then given by [marzetta2016fundamentals]
@@ IV-A Resource Allocation
Power allocation: We adopt the fractional power allocation method [demir2021foundations], which assigns more power to users with larger values of $\gamma_{mk}$, $m\in\mathcal{A}_{\mathsf{active}}$, where we recall that $\gamma_{mk}$ denotes the estimated channel gain as defined in (3). The power allocated by AP $m$ to user $k$ is then obtained as
@@ IV-B DRL Agent Formulation
With the resource allocation mechanisms described above, antenna selection and power allocation for EE maximization reduce to the joint optimization of $\{\zeta,\kappa,\nu\}$. We formulate this as a reinforcement learning (RL) task [sutton1998reinforcement], where a RL agent adaptively adjusts these parameters based on the observed network states. It is worth noting that all of these parameters depend only on the large-scale fading coefficients $\beta_{mk}$. Consequently, the stage remains unchanged within one large-scale fading interval. Let $t\in\{1,2,\ldots,T\}$ index the large-scale intervals, hereafter referred to as “time slots” for simplicity.
@@ IV-C Proximal Policy Optimization
To solve the RL problem in (16), we employ proximal policy optimization (PPO) [schulman2017proximal], which effectively handles continuous variables and ensures stable learning by restricting policy updates to small, controlled steps. In PPO, the agent’s policy $\pi_{\theta}({\mathbf{a}}_{t}|{\mathbf{s}}_{t})$ and value function $V_{\phi}({\mathbf{s}}_{t})$ are parameterized by neural networks with parameters $\theta$ and $\phi$, respectively. The policy network outputs the mean and variance of a continuous action distribution from which ${\mathbf{a}}_{t}$ is sampled, while the value network estimates the expected return $V_{\phi}({\mathbf{s}}_{t})$ to guide policy updates. The advantage function, estimated via generalized advantage estimation (GAE) [schulman2017proximal], is given by
@@ V Performance Evaluation
We consider a square network area of $1\times 1$ km2, wherein the APs and users are uniformly and independently distributed, with wrap-around at the edges to avoid boundary effects. The large-scale fading coefficient is modeled as $\beta_{mk}=\mathrm{PL}_{mk}\cdot z_{mk}$, where $z_{mk}$ is log-normal shadowing with standard deviation $\sigma_{\mathrm{sh}}=8$ dB. The path loss $\mathrm{PL}_{mk}$ (in dB) follows the three-slope model [ngo2017total]:
@@
The DRL agent is trained using the PPO algorithm [schulman2017proximal]. Both the actor and critic networks consist of two hidden layers with 256 neurons each and ReLU activations. The learning rate and reward penalty coefficient are set to $3\times 10^{-4}$ and $\xi_{\mathrm{pen}}=20$, respectively, and the agent is trained for 300,000 timesteps using a minibatch size of 64. Other key hyperparameters of the PPO algorithm are the discount factor $\vartheta=0.99$, the clipping parameter $\epsilon=0.2$, and the GAE parameter $\lambda=0.95$.
We compare the proposed DRL scheme against three benchmarks: The DRL-AP scheme assumes all APs are active ($\zeta=1$) and optimizes only antenna allocation ($\kappa$) and power control ($\nu$), isolating the effect of AP deactivation. The DRL-AO scheme optimizes only AP activation ($\zeta$), with $\kappa=0$ and $\nu=1$, highlighting the benefits of network sparsification. Finally, FA-SCA uses all antennas and optimizing power allocation using the SCA method [ngo2017total].